OpenStudy (anonymous):

Put the function P = 194(0.2)^t in the form P = P0e^kt. (Round your answers to four decimal places.) P0= k=

OpenStudy (anonymous):

\(\huge P=194(\frac{1}{5})^t=194\cdot 5^{-t} \rightarrow \frac{P}{194}=5^{-t} \) \(\huge ln(\frac{P}{194})=ln(5^{-t}) \) \(\huge ln(\frac{P}{194})=-t \cdot ln5 \) \(\huge e^{ln(\frac{P}{194})}=e^{-t \cdot ln5} \) \(\huge \frac{P}{194}=e^{(-ln5)(t)} \) \(\huge P=\color {red}{194} \cdot e^{(\color {blue}{-ln5})(t)} \) \(\huge P=\color {red}{P_0} \cdot e^{\color {blue}kt} \)

OpenStudy (anonymous):

what is k?

OpenStudy (anonymous):

is that a serious question? match up the red and blue quantities....